Dr. Claus-Peter Wirth Brandenburger Str. 42 D–65582 Diez Germany

                                                         wirth@logic.at
                                                       +49 152 07198176

Dr. Wirth, Brandenburger Str. 42, D–65582 Diez

Mr Chris Jowett (SICSA) Informatics Forum 10 Crichton Street Edinburgh EH8 9AB, Scotland Großbritannien

Diez, July 9, 2010

Receipts for my self-covered costs (travel, accomodation, and subsistence) for my stay as SICSA Distinguished Visitor, April–June, 2010, in Scotland

Dear Mr Jowett,

Please find the following receipts enclosed with this letter, numbered and sorted.

Stamps for letter of receipts, see envelope, 7.25 ,  6.45 £
Ferry Netherlands–UK and back, 454.00 , Receipt 1  404.06 £
Fuel for the travel on the Continent, 206.00 , Receipts 2–4  183.34 £
Victuals for the travel on the Ferry, 83.70 , Receipts 5–10  74.49 £
Fuel for the travel within Scotland, Receipts 11–13  122.00 £
Victuals at Sainsbury’s, Receipts 14–31  422.18 £
Single room, Russell Gardens, Edinburgh, Receipts 32, 33  1240.00 £
Essential Scottish culture, Receipts 34, 35  21.00 £
Prepaid SIM Card for my mobile, essential for traveling alone, Receipts 36, 37  43.00 £
Victuals at Tesco’s Receipts 38–41  54.80 £
Heap of Victuals at various places, Receipts 42-62  140.26 £
Heap of Victuals at various places, Receipts 63–68  106.63 £
Heap of Victuals at various places, Receipts 69–82  226.43 £
Heap of Victuals at various places, Receipts 83–90  104.23 £
Tips in Restaurants without included service, conservative estimate  60.00 £
Lost receipts for victuals conservative estimate  60.00 £
Lost fuel bill, conservative estimate  30.00 £
Total self-covered costs  3299.68 £
Received from SICSA in cash  −2510.00 £
Yet uncovered own expenses  789.68 £

For the currency conversion I took the rate of March 30, 2010, when I had to pay for the ferry: 0.89 \frac{£}{}.

In case there is some money left in my SICSA budget, please send it in EURO to my International Bank Account Number DE33 5105 0015 0630 0141 90, SWIFT-BIC: NASSDE55XXX.

Thank you very much again for you kind and effective support.

Yours sincerely,